OpenStudy (anonymous):

16/(x+2) - 3 = (1/3)(q+1)

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

simplify and find x

OpenStudy (anonymous):

Is that "q" another variable?

OpenStudy (anonymous):

oh no its supposed to be an x

OpenStudy (jhannybean):

\[\frac{16}{x+2} - 3= \frac{1}{3}(x+1)\]

OpenStudy (jhannybean):

First simplify the RHS of the equation

OpenStudy (anonymous):

i keep getting different answers

OpenStudy (jhannybean):

What do you get on the RHS when you distribute \(\frac{1}{3}\)?

OpenStudy (anonymous):

1/3x^2+x2-13?

OpenStudy (anonymous):

well 1/3 x^2 + x - 12.3

OpenStudy (jhannybean):

Not quite, i asked just for the RHS :P Breaking this up into steps and then simplifying. RHS I got \(\frac{1}{3}x+\frac{1}{3}\)

OpenStudy (anonymous):

yes same

OpenStudy (jhannybean):

So our equation becomes \[\frac{16}{x+2}-3=\frac{1}{3}x+\frac{1}{3}\] Now we want to add \(+3\) to both sides of the equation. What does your equation become?

OpenStudy (anonymous):

16/x+2= 1/3x + 10/3

OpenStudy (jhannybean):

Awesome. I got the same.

OpenStudy (jhannybean):

Now rewrite it so you've got \[\frac{16}{x+2}=\frac{x}{3}+\frac{10}{3}\]Notice how the two fractions on the RHS have a common denominator of 3? That means we can combine them as one fraction.

OpenStudy (jhannybean):

So we will have \[\frac{16}{x+2}=\frac{x+10}{3}\]

OpenStudy (jhannybean):

Now what do you get when you cross multiply these two fractions?

OpenStudy (sepeario):

\[48 = (x+2)(x+10)\]

OpenStudy (sepeario):

\[48 = x ^{2} + 12x + 20\]

OpenStudy (jhannybean):

That becomes \[x^2+12x-28 =0\]Just factor this.